# BibAgent: An Agentic Framework for Traceable Miscitation Detection in Scientific Literature

### v1

@@ Abstract
Citations are the bedrock of scientific authority, yet their integrity is compromised by widespread miscitations: ranging from nuanced distortions to fabricated references. Systematic citation verification is currently unfeasible; manual review cannot scale to modern publishing volumes, while existing automated tools are restricted by abstract-only analysis or small-scale, domain-specific datasets in part due to the “paywall barrier” of full-text access. We introduce BibAgent, a scalable, end-to-end agentic framework for automated citation verification. BibAgent integrates retrieval, reasoning, and adaptive evidence aggregation, applying distinct strategies for accessible and paywalled sources. For paywalled references, it leverages a novel Evidence Committee mechanism that infers citation validity via downstream citation consensus. To support systematic evaluation, we contribute a 5-category Miscitation Taxonomy and MisciteBench, a massive cross-disciplinary benchmark comprising 6,350 miscitation samples spanning 254 fields. Our results demonstrate that BibAgent  outperforms state-of-the-art Large Language Model (LLM) baselines in citation verification accuracy and interpretability, providing scalable, transparent detection of citation misalignments across the scientific literature.
@@ 1 Introduction
Citations are central to scientific communication, shaping how knowledge claims are established, credit is assigned, and research is evaluated. As such, citations are widely treated as reliable indicators of conceptual connection and research impact, underpinning formal research evaluations for hiring, promotion, and funding (Waltman, 2016). This system, however, rests on the assumption that citations are accurate. Yet, a growing body of scholarship reveals that citation practices are often inconsistent, flawed, or strategically manipulated (Simkin and Roychowdhury, 2003; Wilhite and Fong, 2012; Greenberg, 2009). This phenomenon, miscitation, occurs when cited sources are distorted, fabricated, or taken out of context. The scale is alarming: studies show high rates of inaccuracy, with one analysis finding 39% of sampled biomedical citations were inaccurate (Sarol et al., 2024), and other manual audits reporting misquotation rates between 10%-25%, depending on the field (De Lacey et al., 1985; Jergas and Baethge, 2015; Rekdal, 2014). This is not just accidental; the high-stakes evaluation system has spurred deliberate manipulation, including coercive citation (Wilhite and Fong, 2012), citation cartels (Secchi, 2022), and black market schemes (Chawla, 2024). This long-standing problem is now exacerbated by the emergence of generative AI. Large Language Models (LLMs) introduce new risks by producing “hallucinated citations”, fabricated or distorted references, and their integration into scholarly workflows threatens widespread, systematic errors (Walters and Wilder, 2023; Liang et al., 2024; Rao et al., 2025). Miscitations have severe consequences: they distort the evidentiary basis of claims, propagate inaccuracies, and erode the reliability of citation-based metrics, thereby undermining the integrity of the scientific record and the fairness of the research ecosystem.
Yet, detecting miscitations is inherently challenging. Even for expert reviewers and editors, verifying citation accuracy requires close reading and detailed familiarity with both the citing and cited texts, including the ability to assess whether the citation faithfully represents the original work in content, context, and intent. Given that the average scientific article cites more than 45 references (Dai et al., 2021), and that such assessments demand considerable expertise and time, manual verification is not scalable. The challenge is further compounded by the fact that a substantial portion of scientific literature is behind paywalls or otherwise inaccessible, making it impossible for reviewers to fully evaluate the cited sources.
Automated approaches based on natural language processing (NLP) offer promise, but existing systems face three fundamental limitations. First, the definition of “miscitation” remains conceptually ambiguous. Traditional NLP approaches have historically oscillated between coarse-grained sentiment analysis (Teufel et al., 2006) and binary fact-checking (Athar and Teufel, 2012). These frameworks ignore nuanced errors, such as scope extrapolation or evidence characterization, in which a citation may be technically “supportive” but methodologically misrepresented. Second, current systems suffer from “data myopia” and an “inaccessibility barrier.” High-performing benchmarks like (Wadden et al., 2020) rely primarily on abstracts, yet the evidence required to identify contradictions often resides in other parts of scientific articles, such as experimental tables or appendices. Furthermore, since much of the scientific record remains behind paywalls, AI systems frequently default to silence or “hallucinate” source content when they cannot access the full text. Finally, the emergence of LLMs has introduced the peril of “sycophancy,” wherein models tend to mirror the user’s claims rather than rigorously challenging them against the evidence (Wei et al., 2024). Collectively, these barriers, compounded by a scarcity of labeled data and disciplinary heterogeneity, have limited research to small, proprietary datasets, leaving the scientific community without a robust, generalizable system for maintaining citation integrity in an era of AI-augmented writing.
To bridge these gaps, we present BibAgent, a comprehensive framework designed to evaluate the alignment between citing claims and cited papers for research integrity. Unlike some prior approaches that treat miscitation detection as a single classification task, BibAgent  recognizes that miscitation detection is a multi-layered investigative process that must adapt to varying levels of data accessibility. For accessible sources (i.e., the full text of a cited paper is accessible), we introduce the Accessible Cited Source Verifier (ACSV), which uses an adaptive multi-stage architecture. It moves from efficient bi-encoder retrieval to deep, Large Language Model (LLM)-driven verification only when ambiguity arises. This design ensures traceability and reduces token consumption by 79.4% without compromising the 100% detection rate.
The most significant departure from existing methodologies lies in our treatment of the “Inaccessibility Barrier” through the Inaccessible Cited Source Verifier (ICSV), in case the full text of a cited paper is inaccessible. Rather than defaulting to silence when faced with paywalled text, ICSV aggregates a “voting committee” of open-access papers that cite the same source. By computing a field-normalized consensus among these downstream citers, BibAgent  reconstructs the content of hidden sources through collective community intelligence. This shift-from direct document inspection to a multi-perspective consensus model-allows for a robust chain of integrity even when the primary record is unavailable.
We further introduce two foundational contributions for rigorous evaluation within and beyond the project. First, we formalize a Unified Taxonomy of Miscitation, categorizing errors into five distinct classes ranging from Content Misrepresentation to Scope Extrapolation. Second, we release MisciteBench, a large-scale, cross-disciplinary benchmark of 6,350 expert-validated samples spanning 254 fields, constructed via a “knowledge-blank” protocol to prevent LLM contamination. Together, these resources provide the first robust, generalizable testbed for automated citation verification.
In summary, our core contributions are:
  • BibAgent, the first end-to-end agentic framework capable of handling miscitations involving both accessible and inaccessible cited sources, offering a robust solution to the paywall problem via a community consensus mechanism.
  • Adaptive verification: a multi-stage “zoom-in” logic balancing efficiency and precision, with explicit reasoning trails.
  • Evaluation infrastructure: a comprehensive 5-category Miscitation Taxonomy and MisciteBench, the largest cross-disciplinary benchmark to date, enabling scalable and reproducible assessment of citation integrity.
2 Taxonomy and MisciteBench Construction

### v2

@@ Abstract
Citations are the bedrock of scientific authority, yet their integrity is often compromised by widespread miscitations—ranging from nuanced distortions to fabricated references. Systematic citation verification is currently unfeasible; manual review cannot scale to modern publishing volumes, while existing automated tools are restricted by abstract-only analysis or small-scale, domain-specific datasets in part due to the “paywall barrier” of full-text access. We introduce BibAgent, a scalable, end-to-end agentic framework for traceable citation verification. BibAgent integrates retrieval, reasoning, and adaptive evidence aggregation, applying distinct strategies for accessible and paywalled sources via leveraging a novel Evidence Committee mechanism that infers citation validity via downstream citation consensus. To support systematic evaluation, we introduce MisciteBench, a massive cross-disciplinary benchmark comprising 6,350 miscitation samples spanning 254 fields. Our experimental results demonstrate that BibAgent  outperforms state-of-the-art Large Language Model (LLM) baselines in citation verification accuracy and interpretability, providing scalable, transparent detection of citation misalignments across the scientific literature.
@@
Figure 1: BibAgent : Traceable Miscitation Detection for Accessible and Paywalled Scientific Literature. End-to-end pipeline: DPCM parses papers and maps in-text citations; CSAC routes each reference to ACSV (full-text evidence localization) or ICSV (Evidence Committee over downstream citers) under paywalls; the system emits per-citation verdicts, calibrated confidence, and taxonomy-aligned error codes (AT/CV/CM/SE/EC) with traceable evidence.
@@ 1 Introduction
Citations are the infrastructure of scientific communication: they determine how claims are justified, how credit is allocated, and how careers and funding decisions are made (Waltman, 2016). In principle, a citation is a verifiable contract between a statement and its evidence. In practice, that contract is frequently broken. Large-scale audits report that 39% of sampled biomedical citations are inaccurate (Sarol et al., 2024), and misquotation rates of 10%–25% have been documented across disciplines (De Lacey et al., 1985; Jergas and Baethge, 2015; Rekdal, 2014). For instance, clinical guidelines sometimes still cite retracted trials as primary evidence, or LLM-written drafts attribute causal claims to purely correlational studies. We use miscitation to denote failures where a cited source is distorted, fabricated, or invoked outside its valid evidential role. This problem is newly urgent in the era of generative AI: Large Language Models (LLMs) can produce fluent prose accompanied by plausible but hallucinated or misused references at scale (Walters and Wilder, 2023; Liang et al., 2024; Rao et al., 2025). As a result, scientific claims increasingly risk being backed by references that are unread, misrepresented, or even non-existent, undermining both the integrity of the scholarly record and the fairness of evaluation.
Manual citation checking does not scale. Even expert reviewers must cross-read long papers, resolve context and intent, and judge subtle distinctions (e.g., association vs. causation, narrow vs. broad scope). With modern papers citing more than 45 references on average (Dai et al., 2021), exhaustive verification at review or publication time is infeasible. This has motivated automated approaches, including citation-function and citation-sentiment classifiers (Teufel et al., 2006; Athar and Teufel, 2012; Kilicoglu et al., 2019) and claim-verification systems that judge whether short natural-language claims are supported by abstracts or retrieved snippets (Wadden et al., 2020; Li et al., 2021; Wadden et al., 2022; Liu et al., 2024). More recent work targets citation errors directly (Sarol et al., 2024; Zhang and Abernethy, 2024). However, despite rapid progress in LLM-based reasoning, the field still lacks a solution that is simultaneously taxonomy-aware (distinguishing how a citation fails), paywall-robust (operating when full text is inaccessible), and auditable (providing traceable evidence for each decision) across disciplines and deployment settings.
A key reason is that citation verification operates in two regimes that most prior systems conflate. In the accessible regime, the bottleneck is not just deciding support vs. refutation, but localizing decisive evidence in long, structured documents and diagnosing fine-grained miscitation types (e.g., scope extrapolation or overstated evidence strength) that cannot be reduced to a binary label. In the inaccessible regime, full text may be behind paywalls or otherwise unavailable; treating these cases as uniform abstention creates a systematic blind spot precisely where verification is most needed. Worse, LLM-based verifiers can exhibit sycophancy (Wei et al., 2024): instead of challenging a citing statement against evidence, they may echo its framing, producing confident but ungrounded justifications. Combined with scarce cross-field labeled data and strong disciplinary heterogeneity, these issues have confined existing methods to narrow domains or unrealistic access assumptions, leaving the community without a general, traceable mechanism for maintaining citation integrity in real-world settings.
We start from the view that miscitation detection is fundamentally an evidence retrieval and reasoning problem under these two regimes, and that a practical verifier must remain informative even when the cited source is not directly readable. Our central idea is therefore to treat paywalls not as a hard failure mode, but as an alternative evidential setting: when a cited paper is inaccessible, a verifier should reconstruct what the research community collectively attributes to it by aggregating downstream citation contexts, while explicitly modeling noise and abstaining when community evidence is too sparse or inconsistent. In both regimes, decisions should be traceable: a reader should be able to see the citing context, the supporting evidence (or reconstructed consensus), and its precise error code.
To operationalize this perspective, we introduce BibAgent, an agentic framework for bibliographic miscitation detection, together with MisciteBench, a cross-disciplinary benchmark and evaluation suite for miscitation. At the core is a unified five-category miscitation taxonomy that serves as both (i) the label space of MisciteBench and (ii) the error-code space predicted by BibAgent. BibAgent is an end-to-end framework that maps in-text citations to bibliography entries, routes each reference to an accessible verification branch (evidence localization and reasoning over full text) or an inaccessible verification branch (community-consensus reconstruction from downstream citers), and outputs a single taxonomy-aligned error code for each detected miscitation with traceable evidence. MisciteBench instantiates the taxonomy at scale as a contamination-controlled benchmark of 6,350 expert-validated miscitation instances spanning all 254 Clarivate Journal Citation Reports (JCR) subject categories (Clarivate, 2025), designed to stress-test citation reasoning under distribution shift rather than reward memorization.
In summary, our contributions are:
  • A unified miscitation taxonomy and benchmark. We propose a five-category miscitation taxonomy that provides an operational error-code space for large, heterogeneous corpora, and instantiate it in MisciteBench, a contamination-controlled, cross-disciplinary benchmark of 6,350 expert-validated miscitation instances.
  • A paywall-robust, auditable verification agent. We introduce BibAgent, an end-to-end bibliographic miscitation detection agent that jointly handles accessible and paywalled sources via specialized verification branches and emits explicit taxonomy-aligned error codes with traceable evidence or reconstructed consensus.
  • Practical miscitation detection at scale. Across MisciteBench and additional corpora, BibAgent improves miscitation detection over strong full-text LLM baselines while substantially reducing computation, and its structured outputs support paper-level citation-integrity summaries with fine-grained, interpretable error codes.
2 Taxonomy & MisciteBench Construction